OpenStudy (anonymous):

Solve. Choose the letter for the best answer. (Must show work!) 1. 1/4s = 4 A. s = 4 B. s = 6 C. s = 16

OpenStudy (shamil98):

Multiply by the reciprocal (4/1) to both sides.

OpenStudy (arbershabani97):

1/4s=4 /x4 1s=16 s=16

OpenStudy (arbershabani97):

C